OpenStudy (anonymous):

Please help!! I am almost done with my entire Algebra!! 9.03] Part 1 − Find the vertex, axis of symmetry, domain, and range of the graph of y = −2x2 − 6x + 1. Show all work for full credit. Part 2 − Using complete sentences, explain how you can determine the axis of symmetry, the domain, and range without graphing y = −2x2 − 6x + 1.

OpenStudy (anonymous):

To start with, we get the vertex form of the equation. y = -2(x^2 + 3x) + 1 -> y = -2(x^2 + 3x + 9/4) + 1 + 2(9/4) y = -2(x + 3/2)^2 + 11/2

OpenStudy (anonymous):

So, this is a parabola opening downward with a maximum at (-3/2, 11/2) which is also the vertex. From this vertex, we also know that the axis of symmetry is at x = -3/2. So, we can put any value we want for x, so the domain is all real numbers. And since the max is at 11/2, the range is y <= 11/2

OpenStudy (anonymous):

Is that it?

OpenStudy (anonymous):

Are there any other questions you have with this problem?

OpenStudy (anonymous):

Thx for ur help!! :D Nope, I got it! Thx! :D

OpenStudy (anonymous):

That's right! you're the smiley-girl!

OpenStudy (anonymous):

Ha! Thx for the nickname! LOl, :)

OpenStudy (anonymous):

uw! Have good turkey!

OpenStudy (anonymous):

You too! :)

OpenStudy (anonymous):